OpenStudy (anonymous):

SOMEONE HELP PLEASE!!! [y-5/y-7]-[y+1/y+7]+[y-35/y^2-49]

OpenStudy (anonymous):

no one helped you with this yet?

OpenStudy (anonymous):

\[\frac{y-5}{y-7}-\frac{y+1}{y+7}+\frac{y-35}{(y-7)(y+7)}\] so you can use a common denominator of \[(y+7)(y-7)\]

OpenStudy (anonymous):

\[\frac{(y-5)(y+7)}{(y-7)(y+7)}-\frac{(y+1)(y-7)}{(y-7)(y+7)}+\frac{y-35}{(y-7)(y+7)}\]

OpenStudy (anonymous):

getting stuck on the site

OpenStudy (anonymous):

you get \[\frac{(y-5)(y+7)-(y+1)(y-7)+y-35}{(y-7)(y+7)}\] \[=\frac{9y-36}{(y-7)(y+7)}\] \[=\frac{9(y-7)}{(y-7)(y+7)}=\frac{9}{y+7}\]

OpenStudy (anonymous):

whew that was exhausting

hero (hero):

satellite, you were supposed to have retired

OpenStudy (anonymous):

lol thanks Sat! lol

OpenStudy (anonymous):

imagine how I feel when I work for about 20mins on a problem and still come up with the wrong solution :(

hero (hero):

Happens even to the best of us